1. If 35ml of 0.15 M HCl was used to neutralize 25 mL of a calcium hydroxide solution, what was the concentration of the base?
1
Expert's answer
2021-05-20T07:54:23-0400

Reaction:

Ca (OH)+ 2HCl --> CaCl+ H2O

V (HCl)=35ml=0.035 L

Mole n (HCl)=Molarity×Volume=

0.15M×0.035L=0.00525mol

n(HCl)=1/2n(Ca (OH)2= 1/2×0.00525mol=0.002625mol

Volume(Ca(OH)2)=25ml=0.025L

Molarity (Ca(OH))=

mole(Ca (OH)2)/Volume (Ca (OH)2)=

0.002625mol/0.025L=0.105 mol/L